OpenStudy (anonymous):

use the properties of exponets to simplify 6*6^3 6^8

OpenStudy (anonymous):

Is that a fraction ?

Parth (parthkohli):

\[\Large \color{MidnightBlue}{\Rightarrow [(6^2)^{36}]^8 = (6^{72})^8 = 6^{576} }\]

OpenStudy (anonymous):

@ParthKohli : From where u got That (36)

OpenStudy (mertsj):

When multiplying, add the exponents provided the bases are the saem.

Parth (parthkohli):

Read the question again

OpenStudy (anonymous):

LOL ,I think i can read ! the question says \[6\times 6^3\] \[6^8\]

OpenStudy (anonymous):

@ParthKohli

OpenStudy (mertsj):

Why don't you just ask 9013955 for clarification?

Parth (parthkohli):

@9013955 ?

OpenStudy (anonymous):

I think he meant to write a fraction so its gonna be 6^1+3-8 =6^-4 = 1/6^4 = 1/1296

OpenStudy (mertsj):

Your guess is as good as any. Anyone else want to guess?